A parallelogram has sides of 20cm and 30cm with an angle of 80° between the two sides. What is the length of each of the diagonals?
Answer by lwsshak3(11628)

A parallelogram has sides of 20cm and 30cm with an angle of 80° between the two sides. What is the length of each of the diagonals?
***
use law of cosines: c^2=a^2+b^2-2abcosx
sum of angles in a parallelogram=360˚
For given problem:
a=20, b=30, x=100˚, c=diagonal
c^2=20^2+30^2-2*20*30cos100˚
=400+900-1200*cos100˚
=1300-(-208.38)=1508.38
c=√1508.38≈38.84
length of each of the diagonals=38.84 cm